OpenStudy (anonymous):

x + 4/3 = 10/7- solve and show work

OpenStudy (anonymous):

First step: Subtract 4/3 from both sides. (10/7 - 4/3). In order to do this you have to find a common denominator (21) so the fractions become 30/21 - 28/21 = 2/11 x = 2/11